### v2

Abstract
We highlight the current state-of-the-art in modeling emulsion rheology, ranging from dilute to jammed dense systems. We focus on analytical and numerical methods developed for calculating, computing, and tracking drop deformation en route to developing constitutive models for flowing emulsions. We identify material properties and dimensionless parameters, collate the small deformation theories and resulting expressions for viscometric quantities, list theoretical and numerical methods, and take stock of challenges for capturing connections between drop deformation, morphology, and rheology of emulsions. We highlight the substantial progress in providing quantitative descriptions of the rheological response using analytical theories, dimensional analysis, and powerful computational fluid dynamics to determine how macroscopic rheological properties emerge from microscopic features, including deformation and dynamics of non-interacting or interacting drops and molecular aspects that control the interfacial properties.
keywords:
[…]
1 Introduction
Emulsions are dispersions of droplets in a suspending continuous liquid phase [1, 2, 3, 4]. Examples of emulsions include food materials like milk, creams, salad dressings, and mayonnaise and cosmetics marketed as lotions and creams. Pharmaceutical formulations like certain eye drops, skin care lotions and oral emulsions are designed such that oil phase serves as a carrier for certain hydrophobic bioactives. Mixing of crude oil and water during petroleum extraction or in oceans after oil spills produces petroleum emulsions. Blends of immiscible polymer solutions or melts that form dispersions of droplets in a suspending liquid are also emulsions. Formulating emulsions with flow properties suitable for processing, applications, and sensory perception involves quests that belong to the realm of rheology, i.e., the science of deformation and flow of simple and complex fluids (or soft matter) [3, 4]. Characterization of emulsion rheology involves the measurement of response to applied stress, strain, or strain rate, typically using specialized equipment called rheometers that are designed to create viscometric flows, or well-defined flow fields [3, 4, 5]. The deformability of drops, the possibility of flow within them, and their coalescence or breakup contribute to flow properties that can be quite distinct from the complex fluids containing dispersed particles, micelles, or macromolecules [1, 2, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Additionally, an emulsion’s stability and flow properties depend on the composition, structure, and mechanical properties of the interface between the dispersed and continuous phases [15, 16, 17, 18, 19]. In this contribution, we highlight how size, shape, concentration, interactions, and interfacial properties of dispersed drops influence droplet concentration-dependent variation in the rheological response of emulsions.
Processing operations like pumping, dispensing, pouring, spreading, and even emulsion stability or shelf-life are influenced by shear viscosity, which characterizes the resistance to shear flows associated with velocity gradients perpendicular to the flow direction [5, 4]. Most published emulsion rheology studies focus on the response to shear flows that commonly arise near solid-liquid interfaces, including flows through channels and around immersed objects and primarily describe magnitude and measurement of viscosity, $\eta$ that quantifies viscous resistance to flow [4, 11, 12, 13, 20, 21, 18]. Stream-wise velocity gradients associated with extensional flows commonly arise in converging channels, porous media, and free-surface flows involving the formation of liquid necks that undergo capillarity-driven pinching [5]. Extensional rheology response profoundly influences processing, applications, and consumer use and perception of emulsions. However, due to longstanding experimental and modeling challenges, studies of extensional rheology of emulsions are less common [22, 23, 24, 25]. Spherical drops deform into ellipsoidal shapes in response to weak velocity gradients [6, 7, 26, 8, 9, 18] and can undergo large deformations in response to strong flows, forming slender bodies and even undergoing capillarity-driven breakup [27, 28, 29, 30]. Emulsification or emulsion formation, mixing and blending liquids, and emulsion rheology are three important class of problems involving drop deformations in response to flow fields [30, 18, 31, 32, 33]. Analytical approaches capture minor or small deformations from spherical shape, but numerical approaches are necessary to model large deformations and breakup or coalescence of drops.
Emulsion drops deformed by velocity gradients display elasticity due to interfacial tension, and after flow stops, drops can recover their unperturbed spherical shape, as it is the minimum energy configuration for fixed drop volume [7]. The characteristic timescale for recovering this interfacial energy-favored state is called relaxation time [7]. The surface tension relaxation time, as it is called sometimes in the emulsion rheology literature, appears in the studies of pinching, coalescence and spreading as viscocapillary time as it captures the interplay of viscous and interfacial stresses [34]. Somewhat analogous elastic response is displayed on stopping of flow by perturbed polymer chains in dilute polymer solution, with a relaxation time providing a measure for the characteristic time over which the entropically favored unperturbed coiled state is recovered [35]. In both dilute emulsions and polymer solutions, this elastic recovery of the unperturbed drop shape or coil configuration is at the heart of viscoelastic behavior, captured as modulus in stress relaxation and oscillatory shear measurements or manifested in rod climbing or steady shear torsional rheometry as non-zero normal stress differences.[4, 5]
In nondilute emulsions and particle suspensions, pairwise and higher order interactions and the local arrangement of discrete drops or particles, referred to as microstructure influence the flow behavior.[4, 5, 11] In-situ visualization or monitoring of microstructural evolution in flow fields by optical or spectroscopic methods shows that the rearrangement of drops and the magnitude of drop deformation and orientation together determine rheological response, including rate variation of shear viscosity and normal stress differences and amplitude and frequency dependent moduli measured using oscillatory shear [18, 30, 27, 28, 36]. In the jammed dense emulsion, the flow behavior is additionally influenced by deformation and flows in interconnected liquid films, leading to a yield stress that must be exceeded before flow can be observed, and typically, viscosity exhibits a deformation rate-dependent or stress-dependent nonlinear response [37, 38, 39, 40].
In this brief review, we highlight theoretical and numerical advances in modeling flows of dilute to dense jammed emulsions. The review is divided into six sections. Section 2 and 3 provide motivation, scope, brief history, definitions, transport equations and dimensional analysis. Section 4 presents small deformation theory and constitutive models for dilute emulsions. Section 5 describes constitutive models and numerical methods developed for non-dilute emulsions. Section 6 is a short survey of jammed dense emulsions, and Section 7 lists a few challenges and opportunities.
2 Classifying emulsions and mapping concentration-dependent rheology
Classifying emulsions Emulsions are classified using many criteria, ranging from the choice of dispersed and suspending liquid, interface composition, application (food, pharmaceutical, personal care and cosmetics, petroleum) and drop size and volume fraction range [1, 2, 3, 41, 42]. Emulsions are often described based on the choice of dispersed and suspending phase, oil-water or water-oil emulsions that can be obtained by mechanical mixing, phase separation, microfluidics, vapor condensation, or biologically, as in milk. Here, oil can refer to vegetable oils, crude oil (or derived oil), silicone oils, polymerizable monomers (in latex), or even organic liquids, while the water phase can be made with an aqueous solution or water-based mixed solvent. Both milk and mayo are examples of oil-water emulsions, containing water as the suspending or continuous liquid. Distinct from such emulsions are water-in-water emulsions spontaneously formed as complex coacervate forms between two oppositely charged polyelectrolytes and phase separates forming emulsions that are unstable and have short shelf-life [43], though recent studies describe attempts to enhance stability against coalescence [44].
Typical household emulsions like milk, mayonnaise, cosmetic lotions and creams, salad dressings, and fabric softeners appear milky due to scattering by drops with sizes greater than the wavelength of visible light (drop sizes, a > 1 micron). These are examples of macroemulsions and, being thermodynamically unstable, have a finite shelf-life that can be enhanced by reducing drop sizes and size dispersity, diminishing density difference, increasing suspending fluid viscosity and manipulating drop-drop interactions.[1, 2, 3, 14] Like macroemulsions, nanoemulsions (sometimes called miniemulsions) are also thermodynamically unstable, but smaller drop sizes ($a$ = 50-500 nm) and tighter size control lead to prolonged kinetic stability [1, 45, 46, 47]. In contrast, microemulsions that have relatively small drop sizes ($a$ = 10-100 nm) are thermodynamically stable and appear transparent. Classification based on interface composition: surfactant, protein, lipid, particles, polymers, or complexes between these, emphasizes the critical role played by adsorbed species and interfacial rheology on influencing flow properties and stability of emulsions [1, 3, 18].
Figure 1: Emulsion rheology and microstructure as a function of disperse-phase volume fraction. Representative curves show the increase in relative viscosity from dilute to highly-concentrated emulsions, and the increase in elastic modulus (dashed line) and yield stress (continuous line) for dense emulsions; $\phi$ is the disperse-phase volume fraction. Both elastic modulus and yield stress are normalized by a characteristic capillary stress $\sigma/a$.
Concentration-dependent regimes: dilute to dense Constitutive equations that model flow properties of emulsions consider the influence of number density, interactions, and deformation of drops [1, 4, 11, 13, 12, 20, 42]. The exhibited rheological behavior is considered as linear response if the measured flow properties (stress, viscosity or modulus) do not depend on the impelling quantities (stress, strain, or strain rate). Dilute solutions exhibit viscosity or resistance to flow that is comparable to suspending fluid, as can be observed for animal milks, which are examples of emulsions with a relatively low $\phi$. In the dilute regime, the macroscopic properties that capture linear viscoelastic response, including $\eta_{0}$ increase linearly with $\phi$. The deformation and hydrodynamics of each drop in dilute emulsion can be considered independently, by neglecting the influence of hydrodynamic and thermodynamic interactions. In semi-dilute solutions, pairwise interactions make relative viscosity exhibit a nonlinear increase with $\phi$. In concentrated emulsions, drops are so closely packed that drop mobility and deformation become highly restricted by caging or surrounding drops. The shear viscosity exhibits non-Newtonian response for non-dilute emulsions, and elastic effects become progressively stronger with increase in $\phi$. The semi-dilute to highly concentrated emulsions contain a progressively higher $\phi$ (or number density of drops), and influence of associative and repulsive inter-drop interactions and microstructure become manifest and measurable [11, 1, 10, 48].
Figure 1 illustrates that four concentration regimes, dilute, semi-dilute, concentrated and highly concentrated emulsions, can be identified by examining the variation in relative viscosity, $\eta_{r}$ on increasing droplet volume fraction, $\phi$. Here $\eta_{r}=\eta/\mu$ representing the emulsion’s zero shear viscosity scaled with suspending fluid viscosity, $\mu$. Viscosity increases with $\phi$ substantially in the highly concentrated regime, qualitatively emulating the behavior of rigid particle suspensions, where viscosity diverges close to maximum volume fraction [49, 50, 14]. Due to deformability of drops, droplet volume fraction can be increased further leading to the jammed dense emulsion regime. As volume fraction of drops lies beyond the maximum packing fraction for spherical or ellipsoidal particles, jammed dense emulsions contain polygonal-shaped drops separated by interconnected liquid films with a foam-like microstructure. Mayonnaise, an egg-based emulsion of vegetable oil droplets suspended in a aqueous medium [25], is an example of jammed dense emulsion containing closely-packed, polygonal drops, with volume fraction of the drop phase between $\sim 65\%-80\%$. Such dense emulsions display yield stress, $\tau_{Y}$, and elastic modulus, $G$, that increases with volume fraction, [1, 4, 11, 20].The variation in yield stress and modulus scaled by capillary pressure is illustrated in the Figure 1 for jammed dense emulsions. Though it is well-established that increasing drop volume fraction leads to a transition from suspension-like to foam-like behavior as shown schematically, for emulsion drops and for deformable particles, the transition region is dependent on many factors including size and shape, size dispersity, interactions, and mechanisms underlying the deformability of the dispersed phase.[11, 51, 52, 1, 14]
Highlights from 90 years of analytical models for emulsion rheology The review encompasses nearly a century of models that rely on small deformation theory, a perturbation calculation for weak deviations about a spherical shape that are apt for dilute solutions [6, 53, 26, 8, 9, 54, 55]. Taylor (1932) first analyzed drop deformation in the presence of flow [6] and described the viscosity of dilute emulsions by generalizing Einstein’s theory (1906) for a suspension of hard spheres.[56, 57] Decades later, Schowalter, Chaffey and Brenner (1968) [26] extended the model to suggest the existence of normal stress components, but their model reveals no viscosity variation due to drop deformation. Frankel and Acrivos (1970) [9], and Barthès-Biesel and Acrivos (1972) [58] developed constitutive equations for dilute emulsions that describes the response to transient flows. Choi and Schowalter (1975) [10] carried out the extension to semi-dilute solutions, whereas Princen and Kiss (1980s) [37] showed the connection between yield stress or elastic modulus and surface tension for dense emulsions and foams. Flumerfelt (1980) first examined the influence of interfacial tension variation as well as dilatational and shear interfacial viscosity on drop deformation in small deformation limit, and later Leal, Stone and coworkers carried out more extensive examination in the limit large deformation, including the influence of surfactants [19, 59, 54, 29, 18, 16]. Barthès-Biesel (1980) began the examination of deformation and rheology of capsules, defined as viscous drops covered with elastic membranes, and showed that the combination of liquid-like interior enclosed within a solid-like shell leads to behaviors that cannot be inferred from suspension of hard spheres or emulsion containing drops with Newtonian interfaces [60, 61, 62, 63]. Oldroyd (1954, 1955) [7, 64] presented the first attempt at describing the rheology of nondilute emulsions by adopting the effective medium theory proposed in 1946 for a dispersion of deformable particles [65]. Oldroyd also introduced a tensorial framework to capture the complex viscoelastic response of emulsions with appropriate attention to frame invariance. Starting with Taylor’s discussion of drop deformation [6] or with Oldroyd’s framework [7, 64], a large number of analytical and continuum models have been emerged, which incorporate the interplay of drop deformation, interactions, breakup and coalescence processes and rely on numerical and computation approaches especially for connecting microstructure and rheology of nondilute and dense emulsions. We provide selective (and incomplete) but pragmatic overview of the theoretical framework necessary for modeling emulsion rheology.
Scope of this review We provide a brief synopsis of the small-deformation theories based on perturbation methods that are used for capturing drop deformation and rheological response of dilute emulsions to viscometric flows. As dilute emulsions contain non-interacting drops, their shear rheology response under weak flows, including shear viscosity, can be computed by recognizing that contributions from each drop, mildly perturbed drop by the imposed flow must be added to those by the suspending fluid [66, 26, 9, 10, 67, 68, 69, 70, 71, 72, 73, 54, 74, 55]. Deviation from small-deformation conditions are captured by numerical simulations. We organize the discussion according to the composition of the droplet interface in clean drops, and surfactant-covered drops modelled as droplets with surface viscosity. Quantitative descriptions of the rheological response for non-dilute emulsions relies on supplementing analytical theories with computational fluid dynamics to determine the contributions from deformation and dynamics of non-interacting or interacting drops and molecular aspects that control the interfacial properties. We tabulate different methods and highlight their key findings. As the macroscopic rheology response of emulsions is often compared with the expectations of constitutive models developed for suspensions of undeformable particles, we include suitable references for completeness [56, 57].
In this opinion, we exclude discussions relevant to emulsification and highly nonlinear flows of emulsions[31, 32, 33]. We also exclude the discussion of emulsions containing viscoelastic interfaces, drops or suspending liquids [75, 76, 77, 73, 18, 1] and we exclude studies on capsule suspensions [62, 63]. We cite a paucity of datasets and the immensity of challenges involved in theoretical and experimental studies of extensional rheology response as a reason for excluding the few published studies, including our own [25, 23, 78, 79]. We do not cover studies on Pickering emulsions, water-in-water emulsions, microemulsions, and nanoemulsions, and recommend some recent reviews [1, 45, 46, 80, 44, 81]. We exclude any discussion of rheometry techniques and measured rheological response of emulsions or of interfaces enriched with adsorbed species but we anticipate the references included can be used as a guidebook for the road not taken [1, 2, 4, 82, 11, 12, 5, 83, 17, 84, 85]. Capillary pressure, interfacial rheology, disjoining pressure (contributed by intermolecular and surface forces), and bulk rheology of two liquids all influence drainage flows in liquid films separating any pair of deformed drops, and though interplay and drainage kinetics affect emulsions stability and rheology, a comprehensive description of these remains an open challenge [1, 11, 48, 86]. Nevertheless, we plan to highlight reviews, monographs, articles, and textbooks that form essential reading for appreciating state-of-the-art understanding and progress in the experimental, theoretical, and computational studies of emulsion rheology [4, 11, 12, 21, 20, 13, 30, 18, 3].
3 Microhydrodynamics of emulsions: the governing equations and dimensional analysis
3.1 Governing equations and boundary conditions
Emulsions are structured two-phase fluids composed of droplets of density $\rho+\Delta\rho$ and viscosity $\lambda\mu$ suspended in a continuous-phase fluid of density $\rho$ and viscosity $\mu$. If both the dispersed and the continuous phase are Newtonian, incompressible fluids, and interface is also Newtonian and slip or dissipation free, the only additional material parameter included is the interfacial tension that depends on the two liquids chosen. Assuming that variations of an emulsion macroscopic flow occur over a characteristic length scale $L$, the linear momentum and mass conservation equations in the continuum limit, and in the absence of body-force torques, are
$Re\left(\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}\right)=\nabla\cdot\mathbf{\Sigma}\,;\quad\nabla\cdot\mathbf{u}=0\,,$ (1)
where $\mathbf{u}$ is the velocity field averaged over a continuum volume of fluid, $\mathbf{\Sigma}$ is volume-averaged stress tensor in the emulsion, and $Re$ is the macroscopic Reynolds number. Although, both suspended- and continuous fluids are typically Newtonian, macroscopic rheological behavior is non-Newtonian (e.g., shear thinning and normal stress differences) due to the interplay of droplet-level deformation and relaxation, interfacial dynamics, and interdrop interactions leading to an anisotropic emulsion microstruture in response to imposed bulk stresses [87].
In most applications where emulsions play a key role, droplet size is within the micron scale or smaller such that the local Reynolds number defined in terms of the local shear rate and particle size is $Re_{local}=Re(a/L)^{2}$ provided that $a/L\ll 1$, where $a$ is the average, undisturbed droplet size. Hence, the dynamics at the droplet level are governed by the low-Reynolds-number flow equations,
$\mu\nabla^{2}\mathbf{u}-\nabla p+\rho\mathbf{g}=0;\quad\nabla\cdot\mathbf{u}=0$ (2)
$\lambda\mu\nabla^{2}\mathbf{u^{\prime}}-\nabla p^{\prime}+(\rho+\Delta\rho)\mathbf{g}=0;\quad\nabla\cdot\mathbf{u^{\prime}}=0$ (3)
where the primes denote quantities associated with the drop phase, $\mathbf{g}$ is the gravitational acceleration, and $p$ is the mechanical pressure. Equations (2)-(3) are valid everywhere expect at the droplet interface denoted by $S$. Often models assume that the suspending liquid is density matched with the droplet or dispersed phase. Boundary conditions encompass, typically, an imposed flow field
$\mathbf{u}\to\mathbf{u}^{\infty}\quad\text{as}\quad|\mathbf{x}|\to\infty\,,$ (4)
where $\mathbf{x}$ is the position vector measured from the droplet center. In a general case where the droplet interface is covered with a slip layer of a macromolecule, the Navier-slip condition is used
$\mathbf{u}-\mathbf{u^{\prime}}=\alpha(\mathbf{I}-\mathbf{n}\mathbf{n})\cdot(\mathbf{T}\cdot\mathbf{n})\quad\text{for}\quad\mathbf{x}_{S}\in S\,,$ (5)
where $\mathbf{T}$ is the local Newtonian stress tensor, $(\mathbf{I}-\mathbf{n}\mathbf{n})\cdot(\mathbf{T}\cdot\mathbf{n})$ is the tangential component of the stress vector $\mathbf{T}\cdot\mathbf{n}$ at the interface, $\mathbf{x}_{S}$ is a point at the droplet surface, and $\alpha$ is a slip coefficient. Generally, the velocity at the interface is continuous and $\alpha=0$. The traction jump at the interface is given by
$[\mathbf{n}\cdot\mathbf{T}]_{S}=\left(2H\sigma+\Delta\rho\mathbf{g}\cdot\mathbf{x}\right)\mathbf{n}-\nabla_{S}\sigma\quad\text{for}\quad\mathbf{x}_{S}\in S\,,$ (6)
where [.$[.]_{S}$ denotes a jump of the bracketed quantity across the interface, $\nabla_{S}=(\mathbf{I}-\mathbf{n}\mathbf{n})\cdot\nabla$ is the surface gradient operator,
$H=\frac{1}{2}\nabla_{S}\cdot\mathbf{n}$ (7)
is the mean curvature, and $\sigma$ is the interfacial tension coefficient which may vary along the droplet interface in response to gradients in temperature or in-homogeneous distribution of surfactant molecules. In such cases, an equation of state and an evolution equation for surfactant concentration are needed for closure [88, 89]. Typically, a relation between surfactant concentration and surface tension coefficient is given by the non-linear Langmuir equation of state [90],
$\sigma(\Gamma)=\sigma_{0}+RT\Gamma_{\infty}\ln\left(1-\frac{\Gamma}{\Gamma_{\infty}}\right)\,,$ (8)
where $\Gamma$ is the surfactant concentration along the interface, $\sigma_{0}$ is the surface tension of the clean (surfactant-free) interface, $R$ is the ideal gas constant, $T$ is the absolute temperature, and $\Gamma_{\infty}$ is the maximum packing concentration of surfactant molecules in a monolayer. In the absence of flow and after surfactant adsorption occurs for a sufficient time, there is an equilibrium surface tension $\sigma_{eq}$ at which the equilibrium surface pressure $\Pi_{eq}=\sigma_{0}-\sigma_{eq}$ is defined for a given equilibrium surfactant concentration, $\Gamma_{eq}$ [91]. The ratio $\Gamma_{eq}/\Gamma_{\infty}$ known as surface coverage indicates the initial fraction of the interface covered with surfactants. Several adsorption isotherms can be used to model surface tension variations in the presence of surfactants. We direct the interested reader to Table 1 of Ref. [90] for a comprehensive list.
In the limit of dilute bulk concentration of surfactants, the adsorption kinetics and bulk surfactant diffusion are slow compared to local-convective-flow time scales, such that the surfactant layer at the interface is approximately insoluble and follows a time-dependent convection-diffusion equation [88],
$\frac{\partial\Gamma}{\partial t}+\nabla_{S}\cdot(\Gamma\mathbf{u}_{S})-D_{S}\nabla^{2}_{S}\Gamma+2H\Gamma(\mathbf{u}\cdot\mathbf{n})=0\,,$ (9)
where $\mathbf{u}_{S}=(\mathbf{I}-\mathbf{n}\mathbf{n})\cdot\mathbf{u}$ is the tangential component of velocity at the interface, and $D_{S}$ is the surfactant interfacial diffusivity. The second term in Eq. (9) represents surface convection, the third indicates surface diffusion, and the last represents surface dilution due to local changes in interfacial area or surface dilatation.
The evolution of the droplet interface is captured by the kinematic boundary condition,
$\frac{d\mathbf{x_{s}}}{dt}=\mathbf{n}(\mathbf{u}\cdot\mathbf{n})\,.$ (10)
3.2 Relevant physicochemical parameters and dimensionless groups
A characteristic length scale for describing deformation, breakup or coalescence of drops, is the underformed drop size, $a$. The capillary relaxation time defined as $\tau_{\sigma}=\mu a/\sigma_{eq}$, provides a possible characteristic time scale. The characteristic time scale for $\lambda\gg 1$ is defined as $\tau_{\sigma}=\lambda\mu a/\sigma_{eq}$ as the larger of the two viscosities determines the time period for shape relaxation.[9, 30] Assuming a neutrally-buoyant drop ($\Delta\rho=0$) in an imposed linear flow field where $\mathbf{u}^{\infty}\sim\mathbf{x}\cdot\nabla\mathbf{u}$, the characteristic time scale for the flow is $\tau_{f}=\dot{\gamma}^{-1}$, where $\dot{\gamma}$ is the magnitude of the local velocity gradient. The $\tau_{\sigma}$ or viscocapillary time captures the time required to traverse a distance comparable to drop size, with an intrinsic capillary velocity, $\sigma_{eq}/\mu$ set by the ratio of two physicochemical parameters or material properties: interfacial tension and the viscosity.[34] The two material quantities can be used for estimating characteristic scale for pressures or stresses, as follows. The ratio $\sigma_{eq}/a$, provides an estimate for capillary stress. A typical timescale for droplet deformation in shear is $\tau_{d}\sim\tau_{f}=\dot{\gamma}^{-1}$. Setting the undeformed drop size, $a$, as the characteristic length scale, a natural choice for the characteristic velocity is $\dot{\gamma}a$ and hence, from Eqs. (2)-(3) the pressures inside and outside of the droplet scale as $\mu\dot{\gamma}$ and $\lambda\mu\dot{\gamma}$, respectively. The choices of characteristic time, length and stress/pressure scales determine the form of dimensionless equations and boundary conditions obtained after a nondimensionalization of Eqs. (2)-(10).
The dimensionless ratio of viscous and capillary stresses, is defined as the capillary number
$Ca=\frac{\mu\dot{\gamma}}{\sigma_{eq}/a}=\frac{\dot{\gamma}a}{\sigma_{eq}/\mu}=\frac{\tau_{\sigma}}{\tau_{d}}\,.$ (11)
Alternatively, $Ca$ equals the ratio of imposed flow velocity, $\dot{\gamma}a$ to intrinsic capillary velocity, $\sigma_{eq}/\mu$. Capillary number can be equivalently written as ratio of capillary relaxation time to deformation time. Since $Ca$ is also a product of relaxation time, $\tau_{\sigma}$ and deformation rate ($\dot{\gamma}$ for shear), it captures the flow strength in a fashion reminiscent of Weissenberg number $Wi=\dot{\gamma}\tau_{1}$ used in polymer rheology, with $\tau_{1}$ representing the longest relaxation time. Thus, $Ca$ captures the relative magnitude of stress, velocity, and flow strength for calibrating the influence of applied flow conditions on drop deformation and dynamics. Again, for $\lambda\gg 1$, the $Ca$ values should be computed by considering $\tau_{\sigma}=\lambda\mu a/\sigma_{eq}$ as the shape relaxation time [9, 30].
Two more dimensionless groups written as the ratio of stresses or pressures. The Bond number, $Bo$ captures the ratio of hydrostatic to capillary pressures, relevant to determining bouyancy-driven motion and gravity-induced drop deformation. surface tension gradients. The the Marangoni number, $Ma$, is a ratio between distorting viscous stresses and restoring Marangoni stresses that arise due to surface tension variation.
$Bo=\frac{\Delta\rho ga}{\sigma_{eq}/a}\,,\quad Ma^{-1}=\frac{\mu\dot{\gamma}}{\Delta\sigma/a}\,,$ (12)
If the origin of the Marangoni stress is a non-uniform surfactant contribution, then the characteristic magnitude of surface-tension variations equals the magnitude of surface compression modulus $\Delta\sigma=-\Gamma_{eq}\left(\partial\sigma/\partial\Gamma\right)_{\Gamma=\Gamma_{eq}}$ that arises from perturbations about the equilibrium surface concentration, $\Gamma_{eq}$ The dimensionless ratio of $\Delta\sigma$ to ${\sigma}_{eq}$ represented by $\beta$ is therefore a surface elasticity parameter [69, 72].
$\quad\beta=\frac{\Delta\sigma}{\sigma_{eq}}=CaMa\,,\quad Pe_{S}=\frac{\dot{\gamma}a^{2}}{D_{S}}\,,$ (13)
Lastly, $Pe_{S}$ is the surface Péclet number denoting the relative balance between surfactant convection and diffusion along the interface. The process of emulsification by mechanical methods can sometimes require evaluation of inertial effects using the characteristic inertial pressure estimated as ${\rho}U^{2}$. For example, Reynolds number, $Re$ and Weber number, $We$ are defined as the ratio of inertial pressure to viscous and capillary stress, respectively [34].
Dissipative effects stemming from the surface viscosities may affect the dynamics of droplets in flows. Here, a balance between bulk viscous stresses and dissipative interfacial stresses are embedded in two dimensionless Boussinesq numbers,
$Bq_{s}=\frac{\mu_{s}}{\mu a}\,,\quad Bq_{d}=\frac{\mu_{d}}{\mu a}$ (14)
for shear surface viscosity, $\mu_{s}$, and dilational viscosities, $\mu_{d}$, respectively. In such cases, the right-hand side of the traction jump boundary condition in Eq. (6) is augmented by an additive interfacial-viscous traction of the form, $\nabla_{S}\cdot\tau_{S}$, obeying the deviatoric part of the Boussinesq-Scriven constitutive law for Newtonian interfaces [85, 92, 93, 17],
$\mathbf{\tau}^{s}=2\mu_{s}\mathbf{E}_{s}+(\mu_{d}-\mu_{s})(\mathbf{I}_{S}:\mathbf{E}_{S})\mathbf{I}_{S}\,,$ (15)
where $\mathbf{E}_{S}=\frac{1}{2}\left[\nabla_{S}\mathbf{u}\cdot\mathbf{I}_{S}+\mathbf{I}_{S}\cdot(\nabla_{S}\mathbf{u})^{T}\right]$ is the surface rate of deformation tensor, and $\mathbf{I}_{S}=\mathbf{I}-\mathbf{n}\mathbf{n}$ is a surface projector tensor. Consistent with the traction jump in Eq. (6), normalizing Eq. (15) by a characteristic surface stress $\mu\dot{\gamma}a$, characteristic length $a$, and velocity $\dot{\gamma}a$ yields the dimensionless Boussinesq numbers in Eq. (14).
Accounting for surface viscosity alters the interfacial force balance (6) and affects the interfacial transport of surface-active entities on complex interfaces. Gradients in surface tension $\nabla_{S}\sigma$ generate Marangoni stresses that act to immobilize surfactant transport, while surface viscous stresses oppose surface velocity gradients, where shear viscosity has a similar role as Marangoni stresses and surface dilatational viscosity reduces local surfactant concentration by diluation effects [94, 95, 55, 96]. Recent works suggest that surface viscosity depends exponentially on surface pressure [97, 98, 99, 100, 90]
$\mu_{i}=\mu_{i,eq}\exp\left(\frac{\Pi-\Pi_{eq}}{\Pi_{c}}\right)\,,$ (16)
where $\Pi=\sigma_{0}-\sigma$ is surface pressure, $i=s,d$ stands for shear and dilatational viscosities, $\mu_{i,eq}$ and $\Pi_{eq}$ are the equilibrium surface viscosity and surface pressure, respectively, and $\Pi_{c}$ is a characteristic scale of surface pressure variations. Positive values of $\Pi_{c}$ indicate $\Pi$-thickening surfactants, while negative values are used for $\Pi$-thinning surfactants. The relation between surfactant transport and surface viscous stresses is given by combining Eq. 8 and 16 yielding surfactant-concentration-dependent Boussinesq numbers,
$Bq_{i}=Bq_{i,eq}\left(\frac{1-\hat{\Gamma}_{eq}}{1-\hat{\Gamma}}\right)^{\beta/\hat{\Pi}_{c}}\,,$ (17)
where $i=s,d$ indicate the type of surface viscosity, $Bq_{i,eq}$ is a reference equilibrium value, $\hat{\Pi}_{c}=\Pi_{c}/\sigma_{eq}$, Γ^=Γ/$\hat{\Gamma}=\Gamma/\Gamma_{\infty}$, $\hat{\Gamma}_{eq}=\Gamma_{eq}/\Gamma_{\infty}$, and $\beta$ is the elasticity parameter. Typically, the ratio of dilatational to surface viscosity $\lambda_{ds}$ is used to study the relative importance of both surface viscosities.
Emulsions of droplets with slip-boundaries have been used as model system to probe the rheology of emulsions of immiscible polymer blends, where the slip coefficient is defined by the ratio of the interfacial thickness and some isotropic interfacial viscosity [54, 101, 102]. Non-dimensionalizing Eq. (5) yields a dimensionless slip coefficient $\bar{\alpha}=\alpha/(\mu a)$.
3.3 Emulsion macroscopic stress
The continuum, macroscopic volume-averaged stress in Eq. (2) for a particulate system where both dispersed and suspending fluids are Newtonian is
$\mathbf{\Sigma}=\mathbf{\Sigma}^{0}+\phi\mathbf{\Sigma}^{p}\,,$ (18)
where $\phi$ is the drop-phase volume fraction, $\langle$.$\rangle$ . ⟩ denote the volume-average of the quantity in brackets, $\mathbf{\Sigma}^{0}=-\langle p\rangle\mathbf{I}+2\mu\langle\mathbf{E}\rangle$ is the Newtonian stress contribution from the continuous phase, and
$\mathbf{\Sigma}^{p}=\frac{3}{4\pi a^{3}}\frac{1}{N}\sum_{\alpha=1}^{N}\mathbf{S}^{\alpha}\,,$ (19)
is the particle extra stress in a suspension; the sum accounts for the stress contribution of each one of the $N$ particles in suspension given by
[…]
known as the Landau-Batchelor tensor [103], which depends on the surface traction and the velocity distribution over the particle surface, where local low-Reynolds number conditions hold and no external torques are applied. In the limit of a sharp, fluid interface, $\mathbf{\Sigma}\cdot\mathbf{n}\to\Delta\mathbf{f}$ captures the stress jump across the interface defined in Eq. (6). For example, considering clean, neutrally buoyant droplets, $\Delta\mathbf{f}=2H\sigma\mathbf{n}$.
The emulsion shear rheology is defined by a shear viscosity $\Sigma_{12}$ and first- and second-normal stress differences that arise from contributions of the dispersed phase only, [104]
$N_{1}=\phi N^{p}_{1}=\phi(\Sigma^{p}_{11}-\Sigma^{p}_{22})\,,$ (21)
$N_{2}=\phi N^{p}_{2}=\phi(\Sigma^{p}_{22}-\Sigma^{p}_{33})\,.$ (22)
An effective viscosity is derived from the dimensionless form of Eq. (19),
$\eta_{r}=1+\phi Ca^{-1}\bar{\mathbf{\Sigma}}^{p}_{12}\,,$ (23)
where
$\bar{\mathbf{\Sigma}}=\frac{\mathbf{\Sigma}}{\mu\dot{\gamma}}\,,\quad\bar{\mathbf{\Sigma}}^{0}=\frac{\mathbf{\Sigma}^{0}}{\mu\dot{\gamma}}\,,\quad\bar{\mathbf{\Sigma}}^{p}=\frac{\mathbf{\Sigma}^{p}}{\sigma_{eq}/a}\,,$ (24)
and $\eta_{r}\equiv(\eta/\mu)=\bar{\Sigma}_{12}$, where $\eta$ is the emulsion continuum viscosity. Even though both drop and suspending fluids are assumed Newtonian, results from small deformation theories and numerical simulations show that emulsions display a significant shear-thinning behavior and finite normal stress differences ($N_{1}>0$ and $N_{2}<0$) indicative of a characteristic non-Newtonian behavior arising from a balance between the material structure relaxation time $\sim\mu a/\sigma_{eq}$ and imposed flow rates $\dot{\gamma}^{-1}$ reminiscent of the Weissenberg number in elastic soft materials. Emulsion drop polydispersity can be included in the derivation of Eq. (18) if the distribution of particle sizes is known. Equations (18)-(23) hold for the analysis of dilute to concentrated suspensions. At higher concentrations, near the maximum volume fraction of drops, more elaborate constitutive equations are needed to adequately probe the material rheology. The state of stress and flow behavior of jammed dense emulsions are discussed in Section 6.
4 Dilute emulsions: small deformation theory and constitutive models
In this section, we summarize key features of theoretical and numerical investigations of single-drop dynamics and rheology of dilute emulsions by including three cases: clean drops, surfactant-covered drops, and drops with slip at interfaces. We revisit significant theoretical advances made analytically in the two asymptotic limit of small or large droplet deformations in viscometric flows [30, 105]. We mention numerical studies used for bridging the gap between the two asymptotic limits for clean drops [106, 107, 68], surfactant-covered droplets [59, 108, 89, 109, 110, 111, 112, 113, 114], and drops with viscous interfaces [94, 95, 115, 116, 96, 117]. The approaches discussed here form the starting point for investigations on emulsions containing interfaces with non-Newtonian interfacial rheology. For example, proteins or particles as emulsifiers lead to interfacial viscoelasticity or interfacial yield stress and presence of lipid membranes and protein gel networks at interface creates bending modulus, manifested in suspensions of vesicles and cells including blood. We recommend recent reviews and papers for discussions of emulsions containing complex interfaces that exhibit non-Newtonian interfacial rheology.
4.1 Small deformation theories
Taylor’s deformation parameter In 1932, Taylor generalized Einstein’s formula for viscosity a dilute suspension of hard spheres to derive an expression for the viscosity of dilute emulsions in the limit of low $Ca$, clean interface, and for cases with Newtonian dispersed and suspending fluids. Taylor [6, 118] was the first to theoretically and experimentally study the deformation of a neutrally buoyant viscous drop in response to imposed shear or extensional flows, and describe how bulk rheology is informed by drop deformation and orientation at the microscopic scale. For a weakly perturbed spherical drop, the shape change can be measured using a scalar quantity called Taylor’s deformation parameter defined as
$D_{T}=\frac{L-B}{L+B}\,,$ (25)
where $L$ and $B$ are the major and minor axes of the ellipsoid projected onto the velocity-shear rate plane, as shown in Figure 2. In flows with a rotational component of velocity including viscometric shear flows, the ellipsoid (projection of the deformed drop) orients forming inclination angle, $\theta$ measured between the major axis of deformation and the flow direction. For large deformations, especially those encountered in response to extensional flows, $L/B$ is usually used instead of $D_{T}$ [105].
Figure 2: Representative drop deformation in shear and extensional flows; unperturbed shape added as a reference including interfacial properties.
In dilute emulsions, the flow-induced droplet dynamics depend on the physicochemical properties of the two liquids (density and viscosity), composition-dependent properties of the interface (interfacial tension, interfacial rheology, and surface forces), and the strength and type of imposed flow fields (shear and extensional). Qualitatively, the extend of drop deformation and orientation for clean droplets is influenced by an interplay of viscous and capillary stresses dependent on $Ca$, defined appropriately by accounting for interfacial tension, deformation rate, and viscosity ratio $\lambda$ ranging 0 to $\infty$. Droplets may attain steady shapes or undergo transient flow-induced deformation, possibly leading to interfacial instabilities and breakup (e.g., tip-streaming, burst, and thread breakup by Rayleigh instabilities) [27, 119, 30, 105]. We recommend the classical papers by Leal’s group for a comprehensive survey of clean droplet dynamics in unbounded shear and extensional flows [29, 28], and direct interested readers to Guido’s review on droplet deformation in confined flows and viscoelastic fluids [77].
Clean droplet dynamics in unbounded shear flows In weak flows, $Ca\ll 1$, steady shapes are nearly spherical and the inclination angle $\theta\sim 45^{\circ}$, to leading order in $Ca$, as sketched in Figure 2. At higher flow strengths, for a given $\lambda$, droplet shapes become more elongated as $Ca$ increases and the major axis of deformation aligns with the flow direction as the droplet rotates in response to the local vorticity of the flow. In this limit, drops with viscosities below a critical value $\lambda_{c}\sim 4$ may undergo breakup at a critical flow strength $Ca_{c}$, whereas high-viscosity drops remain stable for $\lambda>\lambda_{c}$, for arbitrary $Ca$ [30, 105]. For example, clean droplets with the same viscosity as the suspending medium undergo breakup at a critical value $Ca_{c}\approx 0.43$ [120].
Experiments by Mason and coworkers [121, 27] characterized the drop deformation and breakup modes of clean droplets in shear flow for $\lambda<\lambda_{c}$ and some cases reproduced are in Fig.3(a). The breakup modes depend on a balance between the rate of increase of capillary number up to and across $Ca_{c}$ and droplet relaxation time. For $\lambda<0.2$ and high $dCa/dt$ rates, the droplets experience tip-streaming breakup mode; whereas for low enough $dCa/dt$, tip-streaming breakup may be suppressed and the droplet deforms into a thin-liquid thread and breakup into smaller droplets by Rayleigh instability. However, numerical and experimental results in extensional flows support the assumption that tip-streaming instabilities occur only in the presence of surfactants [29, 122, 109, 96]. Theoretical and numerical analysis on tip-streaming breakup instability remains an active area of research.
In weak extensional flows, clean droplets attain a stable, stationary shape for all $\lambda$, where the droplet principal axis of deformation is aligned with the flow direction of maximum extension, as illustrated in Fig. 3(b). Here, the transient approach to steady shapes is monotonic, since the flow is free of vorticity. For $Ca=O(1)$, two main regimes of droplet steady deformation are of interest: (i) nearly ellipsoidal shapes are observed for moderate and large $\lambda$, (ii) for $\lambda\lesssim 0.1$, droplets deform into shapes with nearly-pointed ends. For larger values of $Ca$, high-viscosity drops deform into slender threads that eventually breakup into smaller droplets. Low-viscosity drops are able to sustain highly elongated shapes for even larger flow strengths, but will breakup into small droplets via Rayleigh-Plateau instability if $Ca\gg Ca_{c}$. Drop relaxation after the flow field is switched off may also lead to drop breakup into a chain of droplets of uniform size if the droplet initial shape is sufficiently elongated by the flow.
$(a)$$(b)$
Figure 3: Schematic diagram of droplet deformation in shear and extensional flows. Image adapted from Ref. [121] for shear flow experiments (a) and from Fig. 9 in Ref. [108] for numerical results in extensional flows (b). The two sets in extensional flows depict snapshots of drop relaxation of clean and surfactant-covered droplets at different dimensionless times as indicated. Details on the experimental data sets in part (a) are listed in Appendix B.
The presence of surface inclusions (e.g., surfactant molecules, proteins, lipids) alters the classical dynamics of transient and steady shapes of clean droplets [18]. For surfactant-covered drops, deviations from the clean droplet deformation are governed by a balance among (i) interface convection of surfactants towards regions of high curvature and stagnation points lowering surface tension locally, (ii) local surfactant dilution due to drop deformation and creation of surface area, and (iii) diffusion of surfactant which tends to homogenize the surfactant distribution along the interface. Gradients in surface tension induce Marangoni stresses which act against surface deformation [19, 59, 55]. The critical $Ca_{c}$ for the onset of unsteady deformation and breakup is usually larger compared to clean droplet results, but it can be smaller depending on flow strength and on the local vorticity of the flow [69]. Figure 3(b) shows the relaxation of clean and surfactant-covered droplets at different times after being initially deformed by an extensional flow. Surfactant redistribution along the droplet surface stabilizes the shape against transient configurations that may lead to droplet breakup.
The qualitative behavior of droplets with viscous interfaces in linear flows introduces an additional surface viscous stress to the force balance Eq. (6), where droplet shape and rheology depend on flow type and emulsion’s composition, for example, the relative contribution of shear and dilatational surface viscosities and their relation to surface pressure and surface tension [115, 95, 55, 117].
4.2 Constitutive models for dilute emulsions
In the limit when a suspended neutrally bouyant, clean droplet deviates from sphericity only slightly, the droplet surface is given by [9, 8, 78]
$S(t)=r(t)-a\left(1+\epsilon\frac{\mathbf{x}\cdot\mathbf{A}(t)\cdot\mathbf{x}}{r^{2}}\right)+O(\epsilon^{2})=0$ (26)
where $\epsilon\ll 1$ is a perturbation parameter, $\mathbf{A}$ is the shape distortion tensor, $a$ is the radius of the undeformed, spherical droplet, and $r=(\mathbf{x}\cdot\mathbf{x})^{1/2}$. The rate of change of the droplet shape depends on the kinematics of the imposed flow $\mathbf{u}^{\infty}=(\mathbf{E}+\mathbf{W})\cdot\mathbf{x}$, where $\mathbf{E}$ and $\mathbf{W}$ are the imposed-flow rate-of-strain and vorticity tensors, respectively. The distortion tensor $\mathbf{A}$ can be used to calculate Taylor deformation parameter, inclination angle in shear flows, and define rheological material properties of the fluid. The evolution of the distortion tensor in a reference frame that translates and rotates with the droplet is [30, 74]
$\epsilon\frac{\partial\mathbf{A}}{\partial\bar{t}}-Ca\bar{\mathbf{W}}\cdot\epsilon\mathbf{A}+\epsilon Ca\mathbf{A}\cdot\bar{\mathbf{W}}=Ca\,c_{0}(\lambda)\bar{\mathbf{E}}$
$-c_{1}(\lambda)\epsilon\mathbf{A}+O(\epsilon Ca,\epsilon^{2})\,,$ (27)
where $c_{0}(\lambda)=5/(2\lambda+3)$, $c_{1}(\lambda)=40(\lambda+1)/[(19\lambda+16)(2\lambda+3)]$. Dimensionless quantities are defined as $\bar{t}=t\,/(\mu a/\sigma)$, $\bar{\mathbf{E}}=\mathbf{E}/\dot{\gamma}$,$\bar{\mathbf{W}}=\mathbf{W}/\dot{\gamma}$, and $|\mathbf{A}|=1$. A derivation of Eq. (4.2) is shown in Appendix A for completeness. The equation captures how the rate of change of $\mathbf{A}$ is contributed by two competing terms. The first term distorts away from spherical shape and is linearly dependent on the rate of strain, whereas the second term restores unperturbed shape and depends on $\mathbf{A}$. The neglected terms of $O(\epsilon^{2})$ correspond to harmonics higher than second, wheras of $O(\epsilon Ca)$ arise from the straining flow acting on the distorted shape [30].
The form of Eq. (4.2) reveals two small deformation regimes: (i) for weak flows (i.e., $\epsilon\sim Ca\ll 1$ and $\lambda=O(1)$, the distortion is limited by strong interfacial tension effect, and (ii) large-$\lambda$ and arbitrary $Ca$ but not too large for flows with sufficient vorticity where $\epsilon\sim\lambda^{-1}\ll 1$. For a given flow type and small parameter $\epsilon$, Eq. (4.2) is solved for the distortion tensor $\mathbf{A}$. Here, we summarize up to second-order deformation theories for clean droplet deformation and rheology in viscometric flows and include results for surfactant-covered drops, interfacially viscous drops, and drops with interfacial slip conditions.
Clean droplets in shear flows For a clean droplet in weak shear flows where $\epsilon=Ca\ll 1$ and $\lambda=O(1)$ [118], the deformation parameter shows a linear dependence on $Ca$ to leading order
$D_{T}=\frac{19\lambda+16}{16\lambda+16}Ca+O(Ca^{2})$ (28)
and the inclination angle is [53]
$\theta=\frac{\pi}{4}-\frac{(19\lambda+16)(2\lambda+3)}{80(1+\lambda)}Ca+O(Ca^{2})\,.$ (29)
The other limit when $Ca=O(1)$ and $\epsilon=\lambda^{-1}\ll 1$, the leading order solutions for the Taylor deformation parameter and inclination angle are
$D_{T}=\frac{5}{4}\lambda^{-1}+O(\lambda^{-2})\quad\theta=\frac{10}{19}\frac{\lambda^{-1}}{Ca}+O(\lambda^{-2})\,.$ (30)
Higher-order theories have been developed; for detailed derivation and formulas see Refs. [123, 78, 72, 74, 55].
Small deformation theory reveals the characteristic rheological behavior of dilute emulsions. For clean drops in shear flows in the weak flow limit when $\epsilon=Ca\ll 1$ and arbitrary $\lambda$, a second-order deformation analysis yields [26, 58, 72]
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{5\lambda+2}{2(\lambda+1)}-D_{0}(\lambda)D_{1}(\lambda)Ca^{2}+O(Ca^{3})\,,$ (31)
$\frac{N^{p}_{1}}{\mu\dot{\gamma}}=10D_{0}(\lambda)^{2}Ca\,,$ (32)
$\frac{N^{p}_{2}}{\mu\dot{\gamma}}=-\frac{1}{2}\frac{N^{p}_{1}}{\mu\dot{\gamma}}-Ca\,D_{0}(\lambda)\frac{3(12+9(1+\lambda)-25(1+\lambda)^{2})}{28(1+\lambda)^{2}}$ (33)
where the coefficients $D_{0}$ and $D_{1}$ are listed in Appendix A.1. Equations (31)-(33) reveal the characteristic shear-thinning behavior of emulsion flows with finite positive and negative first and second normal stress differences.
In the limit when $\epsilon=\lambda^{-1}\ll 1$ for arbitrary $\lambda Ca$, Oliveira & da Cunha [74] developed a second-order perturbation theory in powers of $\lambda^{-1}$ and showed that
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\left(\frac{5}{2}-\frac{25}{4\lambda}\right)+\frac{5}{\lambda}\frac{20/19}{\left[(20/19)^{2}+(\lambda Ca)^{2}\right]}\,,$ (34)
$\frac{N^{p}_{1}}{\mu\dot{\gamma}}=\frac{10}{\lambda}\frac{(\lambda\,Ca)^{2}}{\left[(20/19)^{2}+(\lambda Ca)^{2}\right]}\,,\quad\frac{N^{p}_{2}}{\mu\dot{\gamma}}=-\frac{29}{133}\frac{N^{p}_{1}}{\mu\dot{\gamma}}\,.$ (35)
The shear rheology of high-viscosity drops reveals two limits. When $Ca\ll 1$ or weak flows, emulsions of high-viscosity drops behave as Boger fluids with shear rate independent viscosity and vanishing, but finite normal stress differences; a similar behavior is observed for $Ca=O(1)$.
Surfactant-covered drops Vlahovska et al. [72] extended the small-deformation theory for clean droplets to surfactant-covered drops valid for arbitrary viscosity ratios and elasticity parameter. In weak flows, the deformation and inclination angle at leading order are
$D_{T}=\frac{5}{4}Ca+O(Ca^{3})\,,$ (36)
and
$\theta=\frac{\pi}{4}-\left[\frac{(32+23\lambda)\beta+4+\lambda}{48\beta}\right]Ca+O(Ca^{2})\,.$ (37)
In weak flows free of vorticity, the stationary shape and surfactant distribution are independent of viscosity ratio since Marangoni stresses immobilize the droplet interface [108, 72]. The rheological material functions for drops covered with insoluble surfactants in shear flow are
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{5}{2}-D_{2}(\lambda,\beta)Ca^{2}+O(Ca^{3})\,,$ (38)
$\frac{N^{p}_{1}}{\mu\dot{\gamma}}=\frac{5}{2}\frac{4\beta+1}{\beta}Ca\,,\quad\frac{N^{p}_{2}}{\mu\dot{\gamma}}=-\frac{1}{2}\frac{N^{p}_{1}}{\mu\dot{\gamma}}+\frac{75}{28}Ca\,,$ (39)
where the coefficient $D_{2}$ is defined in Appendix A.1 Note that, in the limit of $Ca\to 0$, inserting Eq. (38) into Eq. (23) yields Einstein’s classical result $1+(5/2)\phi$ and emulsion rheology follows the behavior of a suspension of rigid spheres with vanishing normal stress differences.
Recently, Narsimhan [55] developed a higher order small deformation theory for shape and rheology of drops covered with viscous interfaces expanding from previous classical works by Oldroyd [64] and Flumerfelt [19]. To leading order, in the limit as $\epsilon=Ca\ll 1$ and $\lambda$, $Bq_{s}$,$Bq_{d}\sim O(1)$
$D_{T}=\frac{1}{2}\alpha_{0}Ca\,,\quad\alpha_{0}=\frac{1}{8}\frac{19\lambda+16+24Bq_{d}+8Bq_{s}}{\lambda^{*}+1}$ (40)
$\alpha_{0}$ is the Taylor deformation parameter, $\lambda^{*}=\lambda+(6/5)Bq_{d}+(4/5)Bq_{s}$ is a modified viscosity ratio, and the inclination reduces to
$\theta=\frac{\pi}{4}+\frac{Ca}{2}a^{-1}_{D}\,,$ (41)
where $a_{D}(\lambda,Bq_{s},Bq_{d})$ is an expansion coefficient [55] defined in Appendix A.2. The corresponding analytical formulas for shear rheology are
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{5\lambda+2}{2(\lambda+1)}\,,$ (42)
$\frac{N^{p}_{1}}{\mu\dot{\gamma}}=\frac{8}{5}\alpha^{2}_{0}\,,$ (43)
$\frac{N^{p}_{2}}{\mu\dot{\gamma}}=-\frac{1}{2}\frac{N^{p}_{1}}{\mu\dot{\gamma}}+\frac{3\alpha_{0}}{70}\frac{(25\lambda^{*^{2}}+41\lambda+24Bq_{d}+4)}{(\lambda^{*}+1)^{2}}\,,$ (44)
where shear-thinning effects are $O(Ca^{2})$ contributions [55].
In the other small deformation limit when $\epsilon\ll 1$ and $Ca=O(1)$,
$D_{T}=\frac{1}{2}\hat{a}_{E}(1+\hat{a}_{E})+O(\epsilon^{3})\,,\quad\theta=-\frac{1}{2}\frac{\hat{a}_{D}}{Ca}+O(\epsilon^{2})\,,$ (45)
where the small parameter $\epsilon=\lambda^{-1}$ or $Bq_{s}^{-1}$ for $Bq_{s}\sim Bq_{d}$. The form of the coefficients $\hat{a}_{D}$ and $\hat{a}_{E}$ are shown in the Appendix A.2. In this limit, small-deformation theory indicates that the emulsion of either highly viscous internal or surface viscosities behave approximately as rigid spheres with no shear-thinning and no significant elastic effects. This observation is in agreement with the small-deformation theory for high-viscosity drops in weak flows [72, 74].
Droplets with slip at interfaces Ramanchandran & Leal [54] developed a second order small deformation analysis for drops with interfacial slip in weak flows. The model captures the anomalous decrease in relative viscosity measured in emulsions formed by immiscible polymer blends. The viscometric functions in shear flow are
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{5\lambda(2\bar{\alpha}+1)+2}{2\lambda(5\bar{\alpha}+1)+2}+O(Ca^{2})\,,$ (46)
$\frac{N^{p}_{1}}{\mu\dot{\gamma}}=f(\lambda,\bar{\alpha})Ca\,,\quad\frac{N^{p}_{2}}{\mu\dot{\gamma}}=\left[\frac{g(\lambda,\bar{\alpha})}{4}-\frac{f(\lambda,\bar{\alpha})}{2}\right]Ca\,,$ (47)
where the functions $f$ and $g$ are defined in Appendix A.3, for completeness [54]. In extensional uniaxial flow, the theory predicts
$\frac{\tilde{\mu}/\mu-3}{\phi}=\frac{5\lambda(2\bar{\alpha}+1)+2}{2\lambda(5\bar{\alpha}+1)+2}+\frac{g(\lambda,\bar{\alpha})}{4}Ca+O(Ca^{2})\,,$ (48)
where $\tilde{\mu}=3\mu$ is the Trouton viscosity for the pure suspending fluid ($\phi=0$), and
$\frac{\tilde{\mu}}{\mu}=\frac{\Sigma^{p}_{33}-\Sigma^{p}_{11}}{\mu\dot{\gamma}}=\frac{\Sigma^{p}_{33}-\Sigma^{p}_{22}}{\mu\dot{\gamma}}\,,$ (49)
by definition. The effect of interfacial slip on material functions in shear and extensional flows is more pronounced for values of viscosity ratio $\lambda>O(1)$. Slip has a stronger influence in response to extensional flows than shear. The analytical results indicate that slip hinders droplet deformation and decrease effective viscosity of the emulsion. However, quantitative agreement between theory and experiments is not verified even in the limit of infinite slip, suggesting that additional physical mechanisms might contribute to the pronounced viscosity reduction observed in experiments [124].
5 Non-dilute emulsions: constitutive models and numerical methods
5.1 Constitutive models for semi-dilute and concentrated emulsions
Constitutive equations proposed for non-dilute emulsions aim to account for finite effects of drop deformations, interactions, and microstructure with respect to each other at disperse-phase volume fractions typically above $10\%$. Oldroyd [7] used an effective-medium approach to derive an expression for the effective viscosity of semi-dilute emulsions following a perturbation analysis proposed by Frölich & Sack [65] for suspensions of elastic spheres. Finite effects of disperse-phase volume fraction were included using a cell model. The cell model represents a composite system consisting of a particle surrounded by a volume of suspending fluid, beyond which the emulsion (or a suspension) is seen as a continuum material. This condition is enforced by a modified far-field velocity boundary condition for the disturbance flow generated by the particle (that encompass different particle types, e.g., drops, capsules, vesicles, rigid particles, and blood cells). Specifically, Eq. (4) is evaluated at a truncated far field position $b/a\sim\phi^{-1/3}$, where $b$ is the characteristic size of the cell in which pressure and velocity disturbances are evaluated, and $a$ is particle size. Oldroyd’s effective medium analysis results in
$\eta_{r}=1+\phi\frac{5\lambda+2}{2(\lambda+1)}\left(1+\phi\frac{\left(5\lambda+2\right)}{5(\lambda+1)}\right)\,,$ (50)
for the effective viscosity.
Choi & Schowalter [10] proposed an alternative derivation of effective viscosity of nondilute emulsion by expanding on the stress-averaged, small-deformation theories of Frankel & Acrivos [9] and Cox [8], by accounting for interparticle interactions and higher-order effects of disperse-phase volume fraction. In steady shear flow, the Choi & Schowalter’s constitutive equation yields,
$\eta_{r}=1+\phi\frac{5\lambda+2}{2(\lambda+1)}\left(1+\phi\frac{5}{4}\frac{\left(5\lambda+2\right)}{(\lambda+1)}+O(\phi^{5/3})\right)\,,$ (51)
for relative viscosity of emulsions. The expressions for the normal stress differences are given by Eqs. (30) and (31) in Ref. [10]. The agreement between Eq. (51) and experimental data in Fig. 4 shows that the model up is valid up to terms $O(\phi^{2})$ in the semi-dilute regime. Yaron & Gal-Or [125] proposed a similar model considering a free-surface cell approach to account for surfactant effects in the limit of spherical droplets. Later generalizations of Oldroyd and Choi & Schowalter viscosity models were developed to include non-Newtonian effects of the drop and suspending phases [126, 4], though a lot of open questions remain regarding the influence of viscoelasticity of either liquid or the interface.
5.2 Empirical equations
Empirical relations are often used to capture the effective viscosity of emulsions of spherical droplets ($Ca\to 0$) as a function of $\phi$ in analogy with suspensions of rigid spheres. For example, a modification of classical suspension models yields the following equation [125, 10, 67],
$\eta_{r}=\exp\left(\frac{5/2\,\,\phi}{1-\phi/\phi_{m}}\right)^{\alpha}\,,$ (52)
where relative viscosity, $\eta_{r}\equiv\eta/\mu$ is the zero-shear-rate viscosity normalized by the viscosity of the suspending medium, $\alpha=(2/5+\lambda)/(1+\lambda)$. Here, $\phi_{m}$ is the emulsion maximum volume fraction at which the effective viscosity (52) diverges. The value of $\phi_{m}$ decreases with increasing viscosity ratio ranging from $0.63-0.64$ for high-viscosity drops [4]. In the dilute regime, $\phi\ll 1$, Eq. (52) reduces to Taylor’s result (see Eq. (31)). In the limit when $\lambda\to\infty$ and arbitrary concentrations, Eq. (52) recovers a Krieger-Dougherty-like empirical viscosity relation for suspensions of hard spheres [4].
For finite values of viscosity ratio, an alternative Krieger-Dougherty-like viscosity model is [127]
$\eta_{r}\left[\frac{2\eta_{r}+5\lambda}{2+5\lambda}\right]=\left(1-\phi/\phi_{m}\right)^{-2.5\phi_{m}}\,.$ (53)
Predictions for Eq. (53) compared to experimental data are shown in Fig. 4. The inset shows data plotted on a linear-linear axis. The corresponding plot shown using log-log scale helps to emphasize how well Taylor’s pioneering theory [6] captures the rheology of dilute emulsions (details about properties of dispersed and suspending liquids are included in Appendix B). The comparison of theory and experiments reveals that the Choi-Schowalter model [10, 4] captures the non-linearity introduced by drop-drop interactions in nondilute emulsions, but the impact of higher order interactions and microstructure require a careful consideration for $\phi$ > 0.4 or so. For a comprehensive review on the empirical viscosity models for concentrated emulsions see Ref. [128]. For nondilute emulsions, normal stress differences become important and shear-thinning effects are also observed at higher shear rates [20]. We recommend a few recent comprehensive reviews for detailed discussion of microstructure, interactions and rheology of concentrated emulsions. [11, 20, 1]
Figure 4: Comparison of theory and empirical relations for effective viscosity models of dilute and concentrated emulsions, respectively. Taylor’s effective viscosity relation is obtained by inserting Eq. (31) into Eq. (23) (dash-dotted line), Choi & Schowalter is given by Eq. (51) (dashed line), and Eq. (53) is used for the Krieger-Dougherty-like curve (dotted line). See Appendix B for details on the datasets and models used.
5.3 Numerical methods for concentrated emulsions
In this section, we enumerate representative numerical works on modeling semi-dilute to concentrated emulsion flows. We focus on the flow-induced microstructue of deformable drops in unbounded flows. Beyond the dilute regime, pairwise droplet interactions are affected by finite deformation of the drop interface allowing for hydrodynamic diffusion. Droplet deformation in the near contact is the stabilizing mechanism against coalescence in the absence of van der Waals attraction [129]. Scaling analysis for the near-contact motion between two deformable, clean drops within lubrication regime shows a slow algebraic film drainage $h/h_{0}\sim\lambda/(\dot{\gamma}t)$ for $\dot{\gamma}t=O(1)$, where $h$ is the gap between the drops, and $h_{0}$ is a reference, initial gap width. At long times, the internal circulation immobilizes the near-contact motion preventing coalescence [87].
As the disperse-phase volume fraction increases, as illustrated in Fig. 1, many drop interactions become important and an analytical treatment is limited. In this regime, detailed numerical simulations are often used to investigate flow-induced structuring and rheology of concentrated emulsions. The choice of numerical method depends largely on the system parameters (e.g., drop relaxation time, size distribution, and dispersed-phase concentration) and imposed flow conditions. Depending on the type of problem under investigation, for example, whether changes in drop topology or the near contact approach of droplet pairs are of interest, a balance among accuracy, resolution, meshing techniques, and computational cost play a key role in selecting the appropriate numerical method. Complex fluid flows are inherently multiphysics problems governed by phenomena across lengthscales (e.g., from atomistic to continuum descriptions). Continuum numerical approaches for multiphase flows are typically divided in two main categories: interface capturing and interface tracking methods [130, 131].
Interface capturing and tracking methods Interface tracking methods explicitly track marker points on a grid or a mesh that fits the particle interface; classical examples are Boundary Integral Method (BIM) [132] and Immersed Boundary Method (IBM) [133]. Alternatively, interface capturing methods (e.g., Volume of Fluid Method (VoF) [134], Phase Field Method (PFM) [135], and Level Set Method (LSM) [136]) evolve a field variable across the computation domain where the interface is captured implicitly by a specific value of a field variable, for example, the contour of zeroes of the level set function. At continuum scales, where volume-averaged material properties of the fluid are uniform, the interface between two immiscible fluids is often assumed to have zero thickness hence the definition of sharp or dividing interfaces [137]. Interface tracking methods are efficient and accurate in modeling sharp interfaces and are usually the method of choice when physical parameters vary strongly across an interface. However, topological changes (e.g., coalescence and breakup) are challenging and require highly detailed meshing schemes. Interface capturing methods handle topological changes naturally, whereas, interface tracking methods require additional numerical effort. The challenge of using interface capturing methods to model physical systems where material properties are discontinuous across an interface, may be overcome by a hybrid approach of interface capturing methods and immersed interface methods or ghost fluid methods [138, 139, 101].
Particle-based models At mesoscopic length scales bridging the gap between molecular dynamics and continuum simulations, coarse-grained particle-based models (e.g., Dissipative Particle Dynamics [140]) or kinetic-based models (e.g., Lattice Boltzmann Method [141]) are usually employed giving access to additional physics compared to continuum-based approaches such as the Boundary Integral Method or Level Set Method. However, both mesoscopic methods require large computational costs to achieve refined grid resolution typically needed in handling near-contact interactions among suspended particles accurately.
Appendix C includes as Figure 7 a descriptive map of representative interface tracking, interface capturing, and coarse-grained mesoscopic numerical approaches used in modeling multiphase flows. Figure 5 highlights representative numerical results of concentrated to dense emulsions using some of the methods listed in Fig. 7. For a comprehensive review on numerical methods used in modeling interfacial rheology and sharp-interface methods to solve free-surface flows, the reader is directed to Refs. [93] and [137], respectively; and to Ref. [131] for more details other computational methods for multiphase flows.
[…]
Figure 5: Summary of representative numerical works on concentrated, highly concentrated, and dense emulsions. The images are adapted from references using Boundary Integral Method for small-scale [142] (a) and large-scale simulations [143] (b) of clean drops in shear flow; emulsion flow of surfactant-covered droplets in shear flows [144] (c) and through structured domains [145] (d), Lattice Boltzmann Method for flowing emulsions where stabilization against coalescence can be tuned by a repulsive force [146] (e), Lattice Boltzmann Method for jammed, dense emulsions of slightly deformed droplets [147] (f), Volume of Fluid simulations of flowing concentrated emulsions accounting for irreversible topological transitions [148] (g) and [149] (h); and Dissipative Particle Dynamics for concentrated emulsions of droplets in shear flow [150] (i). Details of each method can be found in Fig. 7.
Examples of numerical works on concentrated emulsions Loewenberg & Hinch [142] used boundary integral simulations and presented one of the first attempts to simulate small-scale numerical analysis of concentrated emulsion flows of clean, deformable drops with dispersed-phase volume fraction $\phi\leq 30\%$. The results showed a strong shear-thinning behavior, with large positive first and negative normal stress differences, where typically $|N1|>|N2|$. This rheological response is illustrated by the microstrucuture anisotropy shown in Fig. 5(a) where droplets are more deformed and aligned with the flow direction (left image), whereas in the vorticity direction the drops are closely packed (righ image). Elongation of the droplets in the flow direction promotes large $N_{1}$ and facilitates the motion of drops past each other. This droplet arrangement reduces the collisional cross-section and local viscous dissipation leading to a shear thinning behavior. A similar system of interacting droplets in concentrated emulsions with $\phi<30\%$ has been investigated including inertial effects on the emulsion rheology and flow-induced drop structure [150]. The authors used Dissipative Particle Dynamics method where droplets are stabilized against coalescence by a strong repulsive force as illustrated in Fig. 5(i); breakup events are not considered.
More recent studies address flow-induced structuring and rheology of highly-concentrated emulsions below critical jamming conditions [143, 151]. Zinchenko & Davis [151] used a large-scale boundary integral simulation to probe the rheology of highly-concentrated emulsions in flows with nontrivial kinematics. Large strains were assumed and disperse-phase volume fraction varied in the range $0.550.45<\phi<$ 0.55. The simulations used 400 drops per periodic cell and improved upon earlier works from the same group [152, 143]. A snapshot of a periodic cell is shown in 5(b). The authors propose a five-parameter, generalized Oldroyd model where the variable parameters are determined from viscometric and extensionmetric base flows. For example, shear viscosity, first- and second-normal stress differences are calculated from shear flows, and extensional viscosity and stress cross-difference from extensional flows. Long-time averaged material properties in mixed shear and pure extensional flows retain the qualitative features obtained in small-scale simulations of monodisperse emulsions $\phi\leq 30\%$ [142].
Numerical analysis of drop-scale deformation and bulk rheology beyond the class of clean, deformable droplets have been mostly restricted to dilute to semi-dilute regimes accounting for surfactant-covered drops or drops with surface viscous dissipation [69, 153, 114]. Recently, Zinchenko & Davis [144] extended their numerical scheme for highly concentrated emulsion of clean drops [151] to drops covered with insoluble surfactants [144] in shear and extensional flows. They studied emulsion flows with dispersed-phase volume fractions $0.60.45<\phi<$ 0.6, viscosity ratio $30.25<\lambda<$ 3, and surfactant elasticity $0.20.05<\beta<$ 0.2. Sophisticated meshing schemes needed to capture highly deformed droplets in nearly jammed dense emulsions and numerical resolution of the near contact phenomena of approaching droplets are challenges faced by researchers in this field. A representative snapshot of highly-concentration emulsion of surfactant-covered droplets in shear flow is shown in Fig. 5(c). Figure 5(d) shows BIM simulations a pair of highly-deformable surfactant-covered droplets flowing through a pore geometry; the color gradient along the surface indicates regions of different surfactant concentration.
Influence of drop coalescence and breakup Transient evolution of the emulsion micro-structure in concentrated emulsions including changes in droplet topology (e.g., breakup and coalescence events) remains an open area of research. The critical effect of flow-induced droplet breakup and fragmentation on the microstructure and rheology of emulsions [120, 154, 155], including wall-effects, external force fields, viscoelastic contributions have been reviewed or studied elsewhere [156, 154, 157, 77, 158, 159, 160].
Coalescence and breakup events may coexist in confined emulsion flows leading to non-trivial rheology. For example, shear bands which are regions of high and low droplet concentrations in the vorticity and flow direction, respectively, have been observed in numerical experiments[161, 148]. Figure 5(g) shows a snapshot of the droplet microstructure in VoF simulations adapted from Ref. [148]. Rosti et. al [149] determined the effective viscosity of concentrated emulsions Using a 3-dimensional VoF method for volume fractions in the range $0.310^{-3}<\phi<$ 0.3 and capillary number, $0.30.1<Ca<$ 0.3. Coalescence events lead to a non-monotonic variation of effective viscosity with $\phi$, with a peak around $\phi\approx 0.20$. Representative droplet shape distribution observed in their VoF simulation is shown in Fig. 5(h). Recently, Girotto et al. [146] used mesoscopic Lattice-Boltzmann method to study the evolution of the microstructure of emulsions as the disperse-phase volume fraction increases from semi-dilute to jammed configurations. The authors included coalescence and breakup events and further studied ageing dynamics effects after the flow is stopped. An evolution of the emulsion droplet network as the concentration increases is shown in Fig. 5(e). For a comprehensive review on numerical aspects and recent progress on the modelling of deformable particles in flows using the Lattice-Boltzmann method see Ref. [162]. Peterson et al. [163] proposed a generalized framework model for droplet breakup in dense emulsion flows using a population balance model coupled to droplet shape evolution.
6 Jammed dense emulsions with polygonal drops in a network of films
As the dispersed-phase volume fraction is increased beyond the highly concentrated regime of flowing emulsions discussed in section 5, the corresponding rheology is highly sensitive to the droplet positional structure, drop size, interparticle forces, and polydispersity. In this regime, an emulsion of repulsive droplets (stabilized against coalescence) transitions from an amorphous glass-like behavior for $\phi_{g}\approx 0.58$ to a jammed dense regime at $\phi\approx\phi_{RCP}$ where the microstructure is dense and randomly packed and $\phi_{RCP}\approx 0.64$. In the limit as $\phi\to 1$, the droplet are compressed into polygonal shapes and separated by thin films of continuous phase fluids that intersect at Plateau borders, and thus develop a microstructure or a castle of polyhedral shapes characteristic of dry foams [164, 4, 11, 1, 20, 39]. In this section, we focus on the structure and rheology of jammed dense emulsions where the droplets are densely packed showing a solid-like behavior under weak loading, and a fluid-like behavior beyond an effective yield stress [165, 37].
As discussed in sections 4-5, emulsions under flow show a non-Newtonian, viscoelastic behavior where elastic effects are typically imparted by the disperse-phase relaxation time in response to bulk stresses. Yield stress emulsions show a viscoplastic response to imposed bulk stresses. The Herschel-Bulkley model is often used for capture the flow behavior for a complex fluid that displays a yield stress, and flows with a power law relationship between stress and deformation rate above yield stress. The three parameter model including a power law exponent, $n$, consistency, $K$, and yield stress, $\tau_{Y}$, can be written as
$\tau=\tau_{Y}+K{\dot{\gamma}^{n}}=\tau_{Y}+\tau_{v}{(\dot{\gamma})}\,$ (54)
The viscoplastic behavior may be qualitatively defined by ratio of yield stress and an imposed characteristic stress,
$Bn=\frac{\tau_{Y}}{\tau_{c}}\,,$ (55)
where $Bn$ is the Bingham number, $\tau_{Y}$ is the material effective yield stress, and $\tau_{c}=\mu_{B}U/L$ is the characteristic stress where $\mu_{B}$ is a viscous parameter, $U$ and $L$ are characteristic velocity and length scale, respectively.
Under small strains compared to $\tau_{Y}$, dense emulsions show a jammed, solid-like behavior with elastic modulus given by
$G\approx\frac{\sigma}{a_{32}}\phi^{1/3}(\phi-\phi_{0})\,,$ (56)
where $\sigma$ is interfacial tension coefficient, $a_{32}=3V/A$ is an volume-to-surface-area mean drop radius, and $\phi_{0}\approx 0.71$ is the limiting volume fraction at which the percolation of the droplet network collapses. The rheology of dense emulsions of non-coalescing droplets including typical flow curves and characteristic viscoelastic behavior described by the storage, $G^{\prime}$, and loss moduli, $G^{\prime\prime}$, subject to linear and non-linear viscoelastic flowing regimes has been well documented in reviews and papers[164, 20, 11, 159, 1, 51], where most of the works are experimental. Theory and numerical aspects of the problem remain an active area of research.
The measurement or observation of an apparent yield stress in jammed dense emulsions and suspension of particles with a relatively wide range of interaction is much easier than describing the underlying mechanism involving dynamics of dispersed drops in teh case of emulsions.[14, 166, 167, 168, 169, 170, 52] The collapse of the amorphous glass-like microstructure signals the transition to a fluid-like behavior where a classical empirical model by Princen and Kiss [37]for the yield stress is
$\tau_{Y}=\frac{\sigma}{a_{32}}\phi^{1/3}Y(\phi)\,,$ (57)
and $Y(\phi)$ is an empirical relation showing a logarithmic dependence on $\phi$ [37]. Several models are proposed as detailed in the review by Kim and Mason [11]. Figure 6 illustrates that two empirical models capture the trends observed experimentally for $\phi$ dependent increase modulus and yield stress. Details including the properties of dispersed and suspending fluid, the expression for computing the two quantities and values used for different constants are listed in the Appendix for completeness. For emulsions that display yield stress, recent experiments using gravity-based rheometry show the possibility of measuring both an extensional yield stress and the power law relation between extensional stress and strain rate using analysis of dripping, though challenges remain in quantitatively describing the underlying mechanisms for strong flows where droplet deformability probably plays a role.[79, 168, 23, 25]
Figure 6: Comparison between elastic modulus and yield stress empirical models for jammed dense emulsions. Data sets obtained from Refs. [171] and [172] for elastic modulus and from Refs. [37] and [38] for yield stress. Empirical models for elastic modulus obtained from Refs. [171] and [173], and for yield stress from [38] and [37]. Both elastic modulus and yield stress are normalized by a characteristic capillary stress $\sigma/a$. See Appendix B for details on the datasets and models used.
Denkov and coworkers [174] argued that the second term or the viscous stress contribution, $\tau_{v}(\dot{\gamma})$ for yielded emulsions can be attributed to the energy dissipation in thin films between neighboring drops sliding along each other. Their model anticipates a power law exponent $n$ = 1/2 if disjoining pressure is neglected, and explain why viscous stress and shear viscosity exhibit $Ca^{1/2}$ and $Ca^{-1/2}$ scaling, respectively for flowing emulsions. An extended version of the model suggests $n$ < 1/2 if interfacial dissipation plays a role and $n$ > 1/2 if disjoining pressure exerts an influence. The model appears to capture the diversity in power law exponents observed experimentally in flowing emulsions [175, 174].
Numerical studies of jammed dense emulsions Emulsions display $\phi$ dependent yield stress, and is often used by experimentalists as a model system for investigating rheological response. Numerically modeling jammed dense emulsions proffers a similar opportunity with the advantage that changes in microstructure below and above yield stress in response to applied stress can be visualized and analyzed, as shown in a recent numerical investigation by Negro et al. [147]. The authors numerically investigated in 2D the yield stress and flow behavior of a model emulsion that contains an amorphous deformable, non-coalescing droplets embedded in a Newtonian fluid, as summarized below.
Negro et al. [147] evolved the droplet dynamics using 2D hybrid Lattice-Boltzmann method and computed hydrodynamics by following the evolution of phase field variables and velocity of the suspending fluid using the Cahn-Hilliard equation. The droplets are stabilized against coalescence by a soft repulsion force providing for a weak overlap between droplets and forming a percolated microstructure. The model system of densely packed droplets of conserved area initially lies in an amorphous, immobile glass-like state in response to an external forcing, $f$ or pressure difference in a parabolic flow. When the forcing is greater than a critical value $f_{c}$, the percolated network yields and the microstructure orders along the flow direction. Even for $f<f_{c}$, numerical results indicate the continuous fluid permeates the immobile droplet network and hence the effective viscosity is large but finite. Yielding transition is marked by droplet mean velocity fluctuations and stick-slip fluid motion. An analysis of bidisperse systems of small and large species reveals a similar phase transition occurs for $f>f_{c}$. In this regime, yielding is followed by an ordered microstructure where large species accumulated near the centerline of the pressure-driven flow and small species are marginated, as shown in Fig. 5(f). This behavior is reminiscent of flow-induced structuring in the bulk and near the boundaries of dilute to concentrated suspensions given by a balance among hydrodynamic diffusion, deformation-induced drift velocity, and local velocity gradient fluxes [176, 177, 178, 179, 180, 181, 129, 182, 183, 184, 185].
7 Challenges, opportunities, and prognosis
Over the past century, the progress in describing the physicochemical origins of the flow behavior of emulsions reflects progress in describing soft matter physics, thermodynamics, intermolecular and surface forces, interfacial properties, and drop deformation, breakup and coalescence. Despite the progress, designing more sustainable, cost-efficient, or functional formulations in form of emulsions remains challenging as many fundamental scientific problems arise. The macromcolecular, supramolecular and particulate ingredients can alter the rheology of dispersed or suspending fluids and influence interfacial properties, affecting stability, application and processing of emulsions. The review captures some highlights from the current state-of-the-art in modeling shear rheology of emulsions containing Newtonian drops in Newtonian continuum phase with a Newtonian interface. Making any of the three non-Newtonian introduces conceptual, characterization and modeling challenges. Additional open questions are encountered in the following contexts, where we restrict discussion to theoretical and computational challenges only.
Extensional rheology response requires carefully considering the impact of large changes in drop shape, leading to the possibility of droplet break-up and coalescence and changes in microstructure that can influence flow response [186]. For nondilute emulsions, there is also a pronounced lack of experimental data that can be used for benchmarking theoretical methods. This is partially due to challenging extensional rheology characterization with a lack of techniques to measure extensional viscosity and visualize drop shape and microstructure evolution in response to practically-relevant deformation rates [25, 23, 24].
Influence of non-Newtonian interfacial rheology Connecting the emulsion rheology response to the specific measures of interfacial rheology response in dilatational, shear, elastic, bending and torsion modes remains a challenge that can benefit from combination of modeling and experimental studies[17]. Adsorbed layers of proteins, surfactants, polymers, particles, and lipids can display interfacial properties ranging from mobile to rigid, spanning theories discussed herein to describing drops with clean interfaces to elastic interfaces (like capsules) [17, 16, 1, 62, 63, 18, 55].
Viscoelastic suspending fluid The constitutive models and numerical studies described in this contribution are inadequate for capturing the rheological response of emulsions containing viscoelastic suspending fluids or viscoelastic droplets in a Newtonian suspending medium. The two-way coupling of bulk elastic stresses to the interfacial stress jump across the interface can be highly-nonlinear introducing challenges in modelling multiphase flows containing moving boundaries. The effect of flow-induced cross-stream migration and deformation of droplets or capsules in viscoelastic background fluids on the rheology of dilute to concentrated suspensions remains an open area of research [187, 188, 1, 77].
Bubbly fluids Theoretical and experimental investigations on the transient rheology of bubble suspensions remain an active area of research [189, 190]. In the limit of emulsions containing bubbles as the dispersed phase ($\lambda\to 0$), Rust & Manga [191, 192] compared small and large deformation theories and numerical calculations to experimental results on the shape, deformation, and effective viscosity of surfactant-free bubbly suspensions.
Role of deformation and processing history, including emulsification Changes in drop sizes and size distributions, and microstructure have a direct impact on the overall flow properties of emulsions. Modeling different emulsification methods and carrying out modeling with polydisperse droplets that emulate the formulations used in personal care, food, or industry-grade emulsions requires deeper dive into rheology and thermodynamics of multi-component systems [193, 81, 44, 1, 16, 3, 2, 41].
Yielding and microstructural evolution of jammed dense emulsions Further research is needed to elucidate the effects of changes in local drop size, shape and number density, and topological changes involving interconnected thin films on the bulk rheology of emulsions, especially if disjoining pressure, interfacial rheology and either dispersed or suspending fluids are viscoelastic [168, 79, 25, 167, 4, 86, 52, 39, 11].
We offer this survey of theoretical and numerical modeling of emulsions rheology to the scientific community, with an awareness despite this remarkable progress, many practical problems remain in producing, storing, processing, and designing emulsions. We anticipate that advances in numerical and computational methods, and emergence of exciting problems and consumer/industry driven quests involving food and personal care emulsions made with sustainable ingredients will drive the field in the near future.
Appendix A Small deformation theory: clean drops distortion tensor
In the limit when suspended neutrally bouyant, clean droplet deviates from sphericity only slightly, the droplet surface is given by [78]
$S(t)=r(t)-a\left(1+\epsilon\frac{\mathbf{x}\cdot\mathbf{A}(t)\cdot\mathbf{x}}{r^{2}}\right)+O(\epsilon^{2})=0$ (58)
where $\epsilon\ll 1$, $\mathbf{A}$ is the shape distortion tensor, $a$ is the radius of the undeformed, spherical droplet, and $r=(\mathbf{x}\cdot\mathbf{x})^{1/2}$. Solution to Eqs. (2)-(3) are obtained assuming a spherical shape by, for example, superposition of vector spherical harmonics. To leading order, shape distortions are captured in the definition of the normal vector $\mathbf{n}=\nabla S(t)/|\nabla S(t)|$ such that [74]
$\mathbf{n}=\frac{\mathbf{x}}{r}-2a\,\epsilon\left[\frac{\mathbf{A}\cdot\mathbf{x}}{r^{2}}-\frac{\mathbf{x}\left(\mathbf{x}\cdot\mathbf{A}\cdot\mathbf{x}\right)}{r^{4}}+O(\epsilon^{2})\right]\,,$ (59)
and hence appears in the calculation of the mean curvature, $H$, given by Eq. (7). Enforcing boundary conditions (5) and (6) at the droplet interface, the leading order interfacial velocity reduces to
$\mathbf{u}_{s}=\mathbf{W}\cdot\mathbf{x}+c_{0}(\lambda)\mathbf{E}\cdot\mathbf{x}-\frac{\sigma}{\mu a}c_{1}(\lambda)\,\epsilon\mathbf{A}\cdot\mathbf{x}\,,$ (60)
where $c_{0}(\lambda)=5/(2\lambda+3)$, $c_{1}(\lambda)=40(\lambda+1)/[(19\lambda+16)(2\lambda+3)]$, and $\mathbf{E}$ and $\mathbf{W}$ are the imposed-flow rate-of-strain and vorticity tensors, respectively, i.e., $\mathbf{u}^{\infty}=(\mathbf{E}+\mathbf{W})\cdot\mathbf{x}$. Inserting Eq. (26) into the kinematic boundary condition (10) written in the form $DS(t)/Dt=0$, where $D/Dt=\partial/\partial t+\mathbf{u}\cdot\nabla$ is the material derivative [194], and using the approximation that $D(\mathbf{x}/r)/Dt\approx\mathbf{W}\cdot\mathbf{x}/r$, $Dr/Dt=(\mathbf{x}/r)\cdot\mathbf{u}_{s}$, and that $\mathbf{W}$ is anti-symmetric yields the evolution equation for the distortion tensor [30, 74].
A.1 Second-order deformation theory coefficients: clean droplets
For completeness, the coefficients appearing in Eqs.
(31)-(22) for clean drops are listed below [72],
$D_{0}=\frac{(19\chi-3)}{20\chi}\,,$ (61)
and
$D_{1}=$ $(-3888-27308\chi+231041\chi^{2}-33637\chi^{3}$ (62)
$-189761\chi^{4}+159201\chi^{5})/(35280\chi^{4})\,,$
and in Eq. (38) for droplets covered with insoluble surfactants [72],
$D_{2}=\frac{5}{1176\beta^{2}}[245\chi+98\beta(3+\chi)+\beta^{2}(-1059+1127\chi)]\,,$ (63)
where $\chi=1+\lambda$.
A.2 Droplets with viscous interface
Coefficients needed in the analytical formulas for
droplets covered with viscous interfaces are listed in this appendix, for completeness. A full analysis for small deformation analytical results in shear and extensional flows are listed in Ref. [55]. The coefficient in inclination angle formula Eq. (41) in the limit when the small parameter $\epsilon=Ca$ is
$a_{D}=$ $[-8(6Bq_{d}+4Bq_{s}+5\lambda+5)]/(64Bq_{d}$ (64)
$+48Bq_{s}+89\lambda+46Bq_{d}\lambda+52Bq_{s}\lambda$
$+38\lambda^{2}+32Bq_{d}Bq_{s}+48)\,.$
The coefficients appearing in Eq. (40) in the limit when $\epsilon=\lambda^{-1}$ and $Bq_{s}\sim Bq_{s}=O(1)$,
$\hat{a}_{D}$ $=-\frac{20}{19}\epsilon\,,$ (65)
$\hat{a}_{E}$ $=\frac{5}{2}\epsilon-\epsilon^{2}(\frac{15}{4}+\frac{5}{38}Bq_{d}+\frac{45}{19}q_{s})\,,$ (66)
and when $\lambda=O(1)$ and $\epsilon=Bq_{i}^{-1}$ where $i=s,d$ and $Bq_{s}\sim Bq_{d}$ are
$\hat{a}_{D}$ $=-\epsilon\left(\frac{3}{2}+\frac{Bq_{s}}{Bq_{d}}\right)\,,$ (67)
$\hat{a}_{E}$ $=\frac{5}{4}\epsilon\left(3+\frac{Bq_{s}}{Bq_{d}}\right)-\frac{5}{64}\epsilon^{2}\left[96+69\lambda+(72+63\lambda)\right.$ (68)
×$\left.\times(Bq_{s}/Bq_{d})+(24+26\lambda)(Bq_{s}/Bq_{d})^{2}\right]\,.$
A.3 Droplets with interfacial slip
The coefficients appearing in Eqs. (46)-(47) for the viscometric functions of droplets with slip in shear flow are [54]
$f=\frac{1}{40}\left[\frac{\lambda(80\bar{\alpha}+19)+16}{\lambda(5\bar{\alpha}+1)+1}\right]^{2}\,,$ (69)
and
$g$ $=\left(3\left[\lambda(80\bar{\alpha}+19)+16\right]\left[5\lambda^{2}(20\bar{\alpha}^{2}+4\bar{\alpha}+5)+4\right.\right.$ (70)
$\left.\left.+\lambda(40\bar{\alpha}+41)\right]\right)/\left(140\left[\lambda(5\bar{\alpha}+1)+1\right]^{3}\right)\,,$
where $\bar{\alpha}$ is the dimensionless slip coefficient defined in section 3.2.
Appendix B Data used in Figs. 3,  4 and 6
B.1 Experimental datasets used in Fig. 3
The schematics redrawn and used in Fig. 3 are adapted from the experimental results detailed in Ref. [121].
  • First row: $\lambda=6$. Silicone oil 30,000 (Dow Corning fluid) in 60 cP oxidized castor oil (Pale 4). Interfacial tension 6.0 dyn/cm.
  • Second row: $\lambda=1$. Oxidized castor oil (Pale 4) in 52.6 cP silicone oil 5000 (Dow Corning fluid). Interfacial tension 4.8 dyn/cm.
  • Third row: $\lambda=0.7$. Oxidized castor oil (Pale 4) in 90 cP corn syrup. Interfacial tension 21 dyn/cm.
  • Fourth row: $\lambda=0.0002$. Distilled water in 52.6 cP silicone oil 5000 (Dow Corning fluid). Interfacial tension 38 dyn/cm.
B.2 Figure 4: $\eta_{r}$ vs $\phi$
Datasets:
  • Squares: obtained from Fig. 8 in Ref. [38] for a monodisperse silicon oil-in-water emulsion with SDS concentration of $10$mM, droplet size $a=0.55\mu$m, viscosity of the oil $\lambda\mu=12$ cP, water viscosity $\mu=0.997$ cP, $\lambda=12$, and $\sigma=9.8$ dyn/cm.
  • Circles: obtained from Fig. 8 in Ref. [38] for a monodisperse silicon oil-in-water emulsion with SDS concentration of $10$ mM, droplet size $a=0.20\mu$m, viscosity of the oil $\lambda\mu=12$ cP, water viscosity $\mu=104$ cP, $\lambda=0.12$, and $\sigma=9.8$ dyn/cm.
  • Pentagons: obtained from Fig. 6 set 2 in Ref. [195] for a polydisperse petroleum oil-in-water emulsion with Triton-X-100 concetration of 2.1 wt%. Effective drop radius $a_{32}=9.12\,\mu$m, viscosity of the oil $\lambda\mu=5.52$ cP, water viscosity $\mu=0.997$ cP, $\lambda=5.54$, and $\sigma=1.5$ dyn/cm.
  • Triangles: obtained from Fig. 6 set 2 in Ref. [195] for a polydisperse petroleum oil-in-water emulsion with Triton-X-100 concetration of 2.1 wt%. Effective drop radius $a_{32}=9.12\,\mu$m, viscosity of the oil $\lambda\mu=5.52$ cP, water viscosity $\mu=0.997$ cP, $\lambda=5.54$, and $\sigma=1.5$ dyn/cm.
Models:
  • Taylor [6]: using [195] emulsion of oil and water viscosities $\lambda\mu=5.52$ cP and $\mu=0.997$ cP, respectively, and $\lambda=5.54$.
  • Choi & Schowalter [10]: using [195] emulsion of oil and water viscosities $\lambda\mu=5.52$ cP and $\mu=0.997$ cP, respectively, and $\lambda=5.54$.
  • Krieger-Dougherty-like: using $\phi_{m}=0.74$ according to Ref. [127].
B.3 Figure 6: $\bar{G}$ vs $\phi$
Datasets:
  • Triangles: $E(\phi)$ points extracted from Fig. 6 in Ref. [171], where $E(\phi)=G\sigma/(a_{32}\phi^{1/3})$. Polydisperse paraffin oil-in-water emulsion with 11.6 wtAlipal CD-128, 58% active. Each emulsion has an individual mean diameter and interfacial tension as follows: $a_{32}=8.43-8.92\,\mu$m, $\lambda\mu=49.2$ cP, $\mu=2.22$ cP, $\lambda=22.2$, and $\sigma=6.20-6.86$ dyn/cm.
  • Circles: dataset for particle size $a=0.53\mu$m from Ref. [172] according to Fig. 2b (black down-triangles) in Ref. [11]. Monodisperse silicon oil-in-water emulsion with SDS concentration of C = 10 mM, where $\lambda\mu=110$ cP, $\mu=0.997$ cP, $\lambda=110$, and $\sigma=9.8$ dyn/cm.
Models:
  • Princen & Kiss [37] model plotted in the range $\phi=0.73-1$.
  • Wilking & Mason [173] model plotted in the range $\phi=0.73-1$; assuming $\phi_{m}=0.71$.
B.4 Figure 6: $\bar{\tau}_{Y}$ vs $\phi$
  • Squares: rescaled data from Table 1 of Ref. [37]. Polydisperse paraffin oil-in-water emulsion with 10% Neodol 25-3S + 2% Neodol 25-9. Each emulsion has an individual mean diameter and interfacial tension in the ranges: $a_{32}=5.73-10.2\mu$m, oil viscosity $\lambda\mu=49.2$ cP, water viscosity $\mu=1.53$ cP, $\lambda=32.2$, and $\sigma=4.50-4.92$ dyn/cm.
  • Triangles: replotted from Fig. 4 (circles) in Ref. [38]. Monodisperse silicon oil-in-water emulsion with SDS concentration of $10$ mM, drop size $a=0.25\mu$m, $\lambda\mu=12$ cP, $\mu=104$ cP, $\lambda=0.12$, and interfacial tension $\sigma=9.8$ dyn/cm.
  • Hexagons: replotted from Fig. 4 (squares) in Ref. [38]. Monodisperse silicon oil-in-water emulsion with SDS concentration of $10$ mM, drop size $a=0.53\mu$m, $\lambda\mu=12$ cP, $\mu=104$ cP, $\lambda=0.12$, and interfacial tension $\sigma=9.8$ dyn/cm.
Models:
  • Princen & Kiss (1989) [37]: plotted in the range $\phi=0.646-1$. Scaled with the following parameters: $a_{32}=10.05\mu$m, $\sigma=4.723$ dyn/cm.
  • Mason, Bibette and Weitz (1996) [38]: plotted in the range $\phi=0.646-1$ using the empirical quadratic fit for the scaled yield stress $\tau_{Y}/(\sigma/a)=0.51(\phi_{eff}-\phi_{c})^{2}$ where $\phi_{c}=0.62$.
Appendix C Summary of representative numerical methods for multiphase flows
This Appendix summarizes in Fig. 7 representative numerical methods commonly used in numerical simulations of multiphase flows, especially those involving the dynamics of moving interfaces.
Figure 7: Mapping of representative numerical methods typically used in simulations of concentrated emulsion flows. The three areas (a), (b), and (c) refer to interface tracking, interface capturing, and particle based methods, respectively. First column shows a general description of each method. The last two columns highlight strengths and weaknesses. Abbreviations used: Finite Element Method (FEM), Finite Difference (FD), Immersed Boundary Method (IBM), Computational Fluid Dynamics (CFD), Partial Differencial Equations (PDEs), Reynolds number (Re), and numerical methods as indicated.